H Studen, ollowing the procedure n this expenimet Studied tre response of red food dye to visible . light ard ariyzed aStock solution of red food dye unkooun conventration. A series of clilutions of dilutions of the Stocr solution of red foodd dye of unoun orvensrath on, f series of dilutians of tme stoct solution of red food dye were mode so +hut a Beers law plot could be prepared. The solutions were prepared aS followS Solution 1.: Idrop of stoct. solution 50.0m of H20 uhon 2: 5.0ml. of solution I and 5.0 mL of t2O oluhon : 5.0 mL of solution 2. and 5.0mL of H-o Solutiont.5.0mL of soluton Band 5.0mL of 20 luton 5, 5.0mL0F Soluh on 나 ond 5.0mL of H2O Laculate the conventration of drops in mL- solutons luhon I using eguation w,ard the other using eguution ord intuble (2) guation o conenttuation, drops mlnumer of drops volume of s olution,ml omentration of soutton = to tal volume ofdluted Solution, mL onn oncentration of diluted solution, drops m-1 -2. Tame toole the .1. transmポures for Solutions 2,3.,and 5 at the analyticel wevelength equivalent absorbances and record them in e are qiven, Convert these transmttancs to SolutionIdroncentration. T- 5.5 22.(p unnown 3 Prepare a plot of +he absorbanve of eacn soluti on o0 the ordidate versus the conventration, n drops mLion ne abscissa 나,,Determine the uonuentration of red foddye n the unknown stock Soluion of red food
Answer #1

According to this question we have to find out the concentration of unknown sample.

Concentration drops (ml-1) = no of drops/ volume of solution (ml)

Concentration drops (ml-1) = 1/50 = 0.02 ml-1

Concentration of sol 1 = 0.02 ml-1

Conc of diluted solution, drops (ml-1) = (Vol of solution 1 (ml)) (conc of sol 1,drops (ml-1))/ total vol of diluted sol

Solution 2

Conc of sol 2 = 5 X 0.02 / 10 = 0.01 ml-1

Solution 3

Conc of sol 3 = 5 X 0.01 / 10 = 0.005 ml-1

Solution 4

Conc of sol 4 = 5 X 0.005 / 10 = 0.0025 ml-1

Solution 5

Conc of sol 5 = 5 X 0.0025 / 10 = 0.00125 ml-1

According to Beer's Lambert law:

A = -log T = log (100/%T)

Solution Conc (ml-1) %T Absorbance
1 0.02 5.5 1.26
2 0.01 22.6 0.65
3 0.005 46.2 0.34
4 0.0025 68.4 0.17
5 0/00125 82.8 0.08
Unknown 51.5 0.29

Now we find the concentration of unknown stock solution

y = mx + b

0.29 = 64x + 0.01

0.29 - 0.01 = 64x

0.28 = 64x

x = 0.28 /64

x = 0.0044 ml-1

So the concentration of unknown stock solution is 4.4 X 10-3 ml-1.